Circular No. 7338

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 Y1 (LINEAR) An object with unusual motion and reported as asteroidal by LINEAR

(discovery observation below) has been found to be cometary in appearance following posting on the NEO Confirmation Page. Z. Moravec, Kleť, reports that the object appeared slightly diffuse with a possible coma of diameter ∼ 10′′ on images taken in poor seeing on Dec. 21 and 22. G. Billings, Calgary, AB, reported an apparent nebulosity of diameter ∼ 12′′ on Dec. 23 CCD images taken with a 0.36-m reflector, and he noted a faint tail ∼ 20′′ long in p.a. 70^{o} on Dec. 27. S. Nakano, Sumoto, Japan, reports that H. Abe (Yatsuka, 0.26-m reflector) found the comet to be evidently diffuse, T. Kojima (Chiyoda, 0.25-m reflector) found a 10′′ coma and a short tail toward the northeast, and T. Oribe (Saji Observatory, 1.03-m reflector) found a 20′′ coma and a 30′′ tail in p.a. 60^{o}, all on Dec. 27. A. Nakamura, Kuma, Japan, found coma diameter 0′.35 and a faint tail in p.a. 60^{o} on Dec. 27 (0.60-m reflector).

1999 UT  α_{2000}  δ_{2000}  m_{2}
Dec. 20.22291  3^{h}47^{m}47.^{s}39  +73^{o}27′48′′.0  17.5

The following parabolic orbital elements are taken from MPEC 1999-Y28, where the additional reported astrometry also appears:

T = 2001 Mar. 23.798 TT  ω = 181.^{o}845
Ω = 188.490 }2000.0
q = 3.19981 AU  i = 134.843
1999 /00 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 28 3^{h}20.^{m}6 +72^{o}01.^{′}2 4.589 5.215 124.^{o}9 8.^{o}9 16.0
Jan. 7 2 54.40 +69 43.6 4.586 5.149 120.0  9.5 15.9
17 2 36.87 +67 10.8 4.610 5.083 113.5 10.2 15.9
27 2 26.12 +64 35.2 4.658 5.018 105.9 10.9 15.8
Feb. 6 2 20.39 +62 05.3 4.723 4.952  97.7 11.4 15.8
16 2 18.33 +59 46.5 4.799 4.887  89.3 11.7 15.8
26 2 18.92 +57 41.8 4.881 4.823  80.8 11.7 15.8
Mar. 7 2 21.45 +55 52.4 4.963 4.759  72.5 11.5 15.8
17 2 25.40 +54 18.4 5.039 4.695  64.4 11.0 15.7
1999 December 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green